solve the linear system
x+y=13
x-y=5

Answer by tutorcecilia(2152)

x+y=13
x-y=5
_______ [add the terms]
2x+0y=18
.
2x=18 [solve for the x-term]
2x/2=18/2
x=9 [plug (x=9 back into either equation and solve for the y-term]
.
x+y=13
9+y=13
y=13-9
y=4
y=4
.
Check by plugging the values back into either equation and solve:
x+y=13
9+4=13
13=13 [checks out]